OpenStudy (anonymous):

Find the area of an equilateral triangle (regular 3-gon) with the given measurement. 6-inch apothem A = sq. in.

jhonyy9 (jhonyy9):

cos30=6/l sqrt3 6 ----- = ---- 2 l lenght of side l= 12/sqrt3 = (12sqrt3)/3 = 4sqrt3 A=((4sqrt3)*6)/2 = 12sqrt3 (sq.in.)

OpenStudy (anonymous):

theres 3 number then sqrt 3 in the answer blanks..

jhonyy9 (jhonyy9):

???

OpenStudy (anonymous):

the answer box is like ___sqrt3

jhonyy9 (jhonyy9):

yes like how i have wrote there 12sqrt3

OpenStudy (anonymous):

nope that's wrong. there is 3 numbers before the sqrt therefore 12 is wrong.

OpenStudy (anonymous):